OpenStudy (anonymous):

WILL FAN AND MEDAL Line A has a slope of -2 and passes through point (3, 4). Line B has a slope of 1/2 and passes through the origin. On the coordinate plane, graph the point where line A intersects line B

OpenStudy (anonymous):

Line A: y= ax +b, note that a is the slope a= -2 and we know (3,4); 4= (-2)*3+b ==> b=10 and y=(-2x)+10 Line B: y2=0.5*x+d we know it passes through (0,0); 0=0.5*0+d ==> d= 0 and y=0.5x this equations give y for the input x. So if we do this; y(x)=y2(x), we find the results where y values are the same for the x inputs: (-2x)+10=0.5*x > 10 = 2.5*x >x=4 ==>> only for x value 4 do these lines cross. you can see that: y(4)= -8+10= 2 and y2(4)=0.5*4= 2 both result in (4,2)